# Predicting Steady-State Behavior in Complex Networks with Graph Neural Networks

### v2

@@ III.3 Training and Testing Strategy
The performance of GCN and GAT in identifying various dynamic states in model networks is highly accurate. GCN is particularly effective in distinguishing between strongly localized and delocalized states (Fig. 4), while GAT excels at differentiating weakly localized and delocalized states (Fig. 5(c)). Although trained models reliably predict states in model networks, applying them to real-world data presents challenges due to imbalanced state distribution and limited dataset size in the $r_{1}$ and $r_{3}$ regions (Fig. 6). To assess real-world applicability, we trained the GAT model on real-world data sets and achieved reasonable accuracy on test datasets (Fig. 6(a-c)).
III.4 Insights of Training Process
@@
In the above, $\hat{\bf A}^{(i)}$ is the normalized adjacency matrix, ${\bf H}^{(0,i)}$ is the initial feature matrix, and W is the learnable weight matrix. Further, $\bm{h}^{(1,i)}_{j}=\sigma\biggl{(}\sum_{k=1}^{n_{i}}\hat{A}^{(i)}_{jk}\bm{h}^{(0,i)}_{k}{\bf W}\biggr{)}$ is the feature vector of node $j$ in graph $i$ and $j^{th}$ row of updated feature matrix ${\bf H}^{(1,i)}$ (Example 1). Further, ${\bf W}^{(\text{lin})}$ and $b$ are the learnable weights of the linear layer. Finally, $y^{(i)}$ is the true scalar value for $\mathcal{G}_{i}$ and $\hat{y}^{(i)}$ is the predicted IPR value.
@@
Figure 7: Portray the distribution of weight matrices ($\mathbf{W}^{(0)},\mathbf{W}^{(1)},\mathbf{W}^{(2)}$) entries for the three GCN layers during the training process with cycle and star networks (Fig. 4(a)). We show the weights matrix entries for the first five epochs, where the ‘Initial Weights’ infer the initial weight matrix, and the ‘End Weights’ infer the weight matrix after the fourth epochs. The gray color indicated weight matrices during epochs $1$, $2$, and $3$.
We calculate the gradients for the linear layer. We know that each graph $i$ contributes to the overall loss $\mathcal{L}$. Therefore, we accumulate the gradient contributions from each graph when computing the gradient of the loss with respect to the weight matrix ${\bf W}^{(\text{lin})}$ (Example 2). Thus, to obtain the gradient of the loss with respect to the weights ${\bf W}^{(\text{lin})}$, we apply the chain rule
@@
$\frac{\partial\bm{q}^{(i)}_{j}}{\partial{\bf W}}=\hat{A}^{(i)}_{j}{\bf H}^{(0,i)}$ (24)
where $\hat{A}^{(i)}_{j}$ is the $j^{th}$ row of ${\bf\hat{A}}^{(i)}$. Now, we combine the results of the chain rule and get
$\frac{\partial\bm{h}^{(1,i)}_{j}}{\partial{\bf W}}=\sigma^{\prime}\left(\bm{q}^{(i)}_{j}\right)\cdot\hat{A}^{(i)}_{j}{\bf H}^{(0,i)}$ (25)
@@
$\begin{split}\frac{\partial\mathcal{L}}{\partial{\bf W}}&=\frac{2}{N}\sum_{i=1}^{N}\sum_{j=1}^{n_{i}}(\hat{y}^{(i)}-y^{(i)}){\bf W}^{(\text{lin})}\cdot\frac{1}{n_{i}}\cdot\sigma^{\prime}(\bm{q}^{(i)}_{j})\\ &\cdot(\hat{A}^{(i)}_{j}{\bf H}^{(0,i)})\end{split}$ (26)
@@
We observe the weight matrices of different layers during the training time. The distribution of the weight matrices provides a visual representation of the weights learned during the training process (Fig. 7). The magnitude of each weight indicates the importance of the corresponding feature. The higher absolute values in the weight matrices suggest that the feature significantly impacts the model’s predictions. We can observe that for different layers, weight matrix values are initially spread evenly around from zeros, but as time progresses, values become close to zeros (Fig. 7).
IV Conclusion
Using the graph neural network, we introduce a framework to predict the localized and delocalized states of complex unknown networks. We focus on leveraging the rich information embedded in network structures and extracting relevant features for graph regression. Specifically, a GCN model is employed to predict inverse participation ratio values for unknown networks and consequently identify localized or delocalized states. Our approach provides a graph neural network alternative to the traditional principal eigenvalue analysis [20] for understanding the behavior of linear dynamical processes in real-world systems at steady state. This method offers near real-time insight into the structural properties of underlying networks. A key advantage of the proposed framework is its ability to train on small networks and generalize to larger networks, achieving an accuracy of nearly $\sim 100\%$ with test unseen model networks to understand delocalized or strongly localized states. This makes the model scale-invariant, with the computational cost of state prediction remaining consistent regardless of network size, apart from the cost of reading the network data.
Our trained GNN framework (e.g., GCN and GAT) effectively identifies three different states in unseen test model network data. Moreover, our model accurately identifies the states in the weakly localized regions for the real-world data. However, distinguishing between delocalized weakly localized and strongly localized states associated with real-world graphs poses a significant challenge. It might be due to the imbalance of data points in different states and limited dataset availability. Moving forward, we aim to address these challenges and improve identification accuracy.

### v3

@@ III.3 Training and Testing Strategy
Further, to understand whether the framework is independent of network size, i.e., “scale-invariant", we substantiate a more consistent study to evaluate the model’s performance on larger graphs. We have trained the GAT model with a network size of 5001500500-1500, $1000$ for each network type. The trained models have been tested on a network size of 5000150005000-15000, $200$ for each network type, and use the threshold function (Eq. 10). Fig. 6 represents the confusion matrices corresponding to the binary classification (i.e., delocalized and localized) and three-class classification (i.e., delocalized, weakly localized, and strongly localized) with $95\%$ and $96\%$ accuracy, respectively. Although trained models reliably predict states in model networks, applying them to real-world data presents challenges due to imbalanced state distribution and limited dataset size in the $r_{1}$ and $r_{3}$ regions (Fig. 7). To assess real-world applicability, we trained the GAT model on real-world data sets and achieved reasonable accuracy on test datasets (Fig. 7(a-c)).
Figure 7: Prediction of Graph Attention Networks (GAT) on real world data sets (ENZYMES, $\texttt{NCI}1$ and $\texttt{MCF}-7$) [26, 27]. (a-c) Our model can also predict the IPR value during testing for real-world graph data sets. We can observe the GAT model predicts the state’s IPR value with significant accuracy. (b) We train the model for $200$ epochs for the real-world dataset. We choose $\tau_{1}=0.05$, $\tau_{2}=0.2$, $\epsilon=1e-6$.
The performance of GCN and GAT in identifying various dynamic states in model networks is highly accurate. GCN is particularly effective in distinguishing between strongly localized and delocalized states (Fig. 4), while GAT excels at differentiating weakly localized and delocalized states (Fig. 5(c)). We observed that the GCN model tends to underperform for specific graph topologies, especially for scale-free and ER random graphs, when tasked with predicting the IPR values. One possible reason is that GCN relies on spectral convolution using normalized adjacency information, which assumes smoothness in node features across the graph [30, 31]. ER networks generally lack sufficient structural regularity [31], and scale-free networks exhibit highly skewed degree distributions and strong local heterogeneities. These features can contravene the homophily assumption and reduce the effectiveness of GCN’s neighborhood aggregation. On the contrary, GAT model leverages learnable attention mechanisms that adaptively weight neighbours based on feature relevance rather than uniform or degree-based contributions. This accedes to GAT for better capture of local variations and importance, which could be particularly beneficial for the topologies. While GAT consistently performs better in specific regimes, we prefer to retain both models in our study to provide a comparative understanding of model behavior across diverse graph topologies. It is crucial to highlight that GCN is always faster and easier to understand [32] than other models. This dual-model perspective allows GCN to serve as a valuable diagnostic tool to better isolate the influence of topology on model performance. Nevertheless, preferring GAT as the primary model for deployment or end-use scenarios could be a more reasonable and practical solution when robustness across topologies is essential.
IV Ablation Study
To determine the most relevant node features among degree centrality, clustering coefficient, betweenness, closeness, and PageRank, we trained GAT models with individual features separately and compared their accuracy with test data. We observed that among the five features, the GAT model trained only on betweenness centrality achieved the best prediction accuracy on the test data (Fig. 8(a)), while the model using the clustering coefficient yielded the weakest performance in classifying between localized and delocalized states (Fig. 8(b)). Accuracy is the proportion of correctly predicted classes among the total number of tests made.
Figure 8: Confusion matrices for the GAT model trained with single node-features. (a, b) Binary classification (strongly localized vs. delocalized) results for betweenness and clustering coefficient features, respectively. (c, d) Three-class classification (strongly localized, weakly localized, and delocalized) results for closeness and clustering coefficient features, respectively. Betweenness and closeness centralities yield the highest accuracies ($\sim 98\%$ and $\sim 95.3\%$), while clustering performs the worst in both tasks.
Based on prediction accuracy on the localized and delocalized classification tasks, we can rank the importance of these features as betweenness ($\sim 98\%$), closeness ($\sim 93\%$), PageRank ($\sim 90\%$), degree centrality ($\sim 86\%$), and clustering coefficient ($\sim 66\%$).
Figure 9: Comparison of prediction performance using different subsets of node features. (a–c) show the results for the GAT model trained using only the top two features: betweenness and closeness centralities, demonstrating high accuracy in regression and classification. (d–f) show results using all five features, yielding similar classification accuracy but slightly better regression slope alignment with actual values. (g–i) display performance using only the bottom three features (degree, PageRank, and clustering coefficient), where classification between delocalized and localized is reasonable, but the model fails to accurately predict the weakly localized class.
Further, for the three-class classification in strong localized, weak localized, and delocalized states the GAT model trained with the closeness feature shows the highest accuracy (Fig. 8(c)) and clustering coefficient shows the lowest (Fig. 8(d)). The ranking of the features based on accuracy is as closeness ($\sim 95.3\%$), betweenness ($\sim 94\%$), degree centrality ($\sim 68\%$), PageRank ($\sim 66\%$), and clustering coefficient ($\sim 57\%$).
Combining the accuracy from binary and three-class classification shows that the top two important features are betweenness and closeness, and the cluster coefficient is the least important. To verify the impact of the top two features on the overall prediction accuracy, we have further trained a GAT model with these two features only. Fig. 9(a-c) shows high prediction accuracy with these two features, which indicates that among all five features, these two features are more relevant and one can only choose these two features to design the solution framework.
Figure 10: Prediction performance of the GAT model after hyperparameter optimization using Optuna. (a) represent the predicted versus actual values of the IPR, (b) confusion matrix of the binary classification and (c) represents the confusion matrix of delocalized, weakly localized, and strongly localized states, respectively. The optimized model uses $256$ hidden channels, $(4,2)$ attention heads in the two GAT layers, dropout rate of $0.48$, and learning rate of $3\times 10^{-3}$. (d-f) Illustrates the results for $3$, 55, and $10$-layer GAT, respectively.
Further, we compared the accuracy of this experiment against the trained GAT model with all five features (Fig. 9(d-f)). It shows similar accuracy to that obtained from the top two features. However, the classification accuracy may not increase much with all five features, but the slope of the regression line towards the actual values is slightly improved due to the use of all five features. Therefore, choosing all five features will help to enhance the solution with marginally improved accuracy. Finally, to understand the impact of the bottom three features (i.e., degree centrality, pagerank, and cluster coefficient), we have trained a GAT model with only these three features and observed that it hugely impacts the regression solution (Fig. 9(g-i)) and consequently the three-class classification. The regression values for delocalized and strongly localized are predicted well; hence, the accuracy of binary classification is good; however, it poorly predicts the weakly localized, and as a result, the prediction accuracy of multi-classification is low. Note that the training data set comprised networks with sizes ranging from $200$ to $300$ nodes, with $1000$ samples per network type (cycle, path, wheel, star, ER, and SF), totaling $6000$ networks. The test data set included networks between $400$ and $500$ nodes, with $500$ samples per network type.
Hyperparameter Optimization: We employed Optuna [33], a hyperparameter optimization framework, to systematically fine-tune our GAT model. The optimization has been carried out based on both architectural and training-related hyperparameters. We designed two experimental setup (a) in the first, we fixed the number of layers to $2$ and optimized other hyperparameters (hidden channels, attention heads per GAT layer, dropout rate, and learning rate), (b) in the second, we varied the number of layers from $1$ to $6$, in addition to tuning other parameters.
For the first set-up (2 layers), the best performance is achieved with $256$ hidden channels, $(4,2)$ attention heads, a dropout rate of $\sim 0.48$, and a learning rate of $\sim 3\times 10^{-3}$. The corresponding accuracy is shown in Fig. 10(a-c). Notably, this model only marginally outperformed the baseline configuration with 64 hidden channels and $(4,1)$ attention heads (Fig. 9 (d-f)). In the second set-up (with variable layers), the best configuration included $5$ layers, $256$ hidden channels, $(2,1)$ attention heads, a dropout rate of $0.41$, and a learning rate of $\sim 2\times 10^{-3}$. The results for this model are shown in Fig. 10 (e).
From the above, we observe that increasing the number of layers significantly raises (nearly $\sim 5$ times, with the same training dataset) training time due to the larger number of trainable parameters, and also classification accuracy does not improve substantially with more layers (Figs. 9 (d) and 10 (a, d-f)). Hence, we prefer to use $2$ or $3$ layers to balance performance and efficiency. This reinforces our earlier observation (Fig. 9 (d)) that shallower GAT architectures can be efficient and effective for our work.
IV.1 Insights of Training Process
@@
In the above, $\hat{\bf A}^{(i)}$ is the normalized adjacency matrix, ${\bf H}^{(0,i)}$ is the initial feature matrix, and W is the learnable weight matrix. Further, $\bm{h}^{(1,i)}_{j}=\sigma\biggl{(}\sum_{k=1}^{n_{i}}\hat{a}^{(i)}_{jk}\bm{h}^{(0,i)}_{k}{\bf W}\biggr{)}$ is the feature vector of node $j$ in graph $i$ and $j^{th}$ row of updated feature matrix ${\bf H}^{(1,i)}$ (Appendix B, Example 1). Further, ${\bf W}^{(\text{lin})}$ and $b$ are the learnable weights of the linear layer. Finally, $y^{(i)}$ is the true scalar value for $\mathcal{G}_{i}$ and $\hat{y}^{(i)}$ is the predicted IPR value.
@@
Figure 11: Portray the distribution of weight matrices ($\mathbf{W}^{(0)},\mathbf{W}^{(1)},\mathbf{W}^{(2)}$) entries for the three GCN layers during the training process with cycle and star networks (Fig. 4(a)). We show the weights matrix entries for the first ten epochs. The gray color indicated weight matrices during epochs 292-9.
We calculate the gradients for the linear layer. We know that each graph $i$ contributes to the overall loss $\mathcal{L}$. Therefore, we accumulate the gradient contributions from each graph when computing the gradient of the loss with respect to the weight matrix ${\bf W}^{(\text{lin})}$ (Appendix B, Example 2). Thus, to obtain the gradient of the loss with respect to the weights ${\bf W}^{(\text{lin})}$, we apply the chain rule
@@
$\frac{\partial\bm{q}^{(i)}_{j}}{\partial{\bf W}}=\hat{\bm{A}}^{(i)}_{j}{\bf H}^{(0,i)}$ (25)
where $\hat{\bm{A}}^{(i)}_{j}$ is the $j^{th}$ row of ${\bf\hat{A}}^{(i)}$. Now, we combine the results of the chain rule and get
$\frac{\partial\bm{h}^{(1,i)}_{j}}{\partial{\bf W}}=\sigma^{\prime}\left(\bm{q}^{(i)}_{j}\right)\cdot\hat{\bm{A}}^{(i)}_{j}{\bf H}^{(0,i)}$ (26)
@@
$\begin{split}\frac{\partial\mathcal{L}}{\partial{\bf W}}&=\frac{2}{N}\sum_{i=1}^{N}\sum_{j=1}^{n_{i}}(\hat{y}^{(i)}-y^{(i)}){\bf W}^{(\text{lin})}\cdot\frac{1}{n_{i}}\cdot\sigma^{\prime}(\bm{q}^{(i)}_{j})\\ &\cdot(\hat{\bm{A}}^{(i)}_{j}{\bf H}^{(0,i)})\end{split}$ (27)
@@
We observe the weight matrices of different layers during the training time. The distribution of the weight matrices provides a visual representation of the weights learned during the training process (Fig. 11). The magnitude of each weight indicates the importance of the corresponding feature. The higher absolute values in the weight matrices suggest that the feature significantly impacts the model’s predictions. For an instance, we can observe that for different layers, weight matrix values are initially spread evenly around from zeros, but as time progresses, values become close to zeros (Fig. 11).
V Conclusion
Using the graph neural network, we introduce a framework to predict the localized and delocalized states of complex unknown networks. We focus on leveraging the rich information embedded in network structures and extracting relevant features for graph regression. Specifically, a GCN model is employed to predict inverse participation ratio values for unknown networks and consequently identify localized or delocalized states. Our approach provides a graph neural network alternative to the traditional principal eigenvalue analysis [20] for understanding the behavior of linear dynamical processes in real-world systems at steady state. This method offers near real-time insight into the structural properties of underlying networks. A key advantage of the proposed framework is its ability to train on small networks and generalize to larger networks, achieving an accuracy of nearly $\sim 95\%$ with test unseen model networks to understand delocalized or strongly localized states. This makes the model scale-invariant, with the computational cost of state prediction remaining consistent regardless of network size, apart from the cost of reading the network data.
Our trained GNN framework (e.g., GCN and GAT) effectively identifies three different states in unseen test model network data. Moreover, our model accurately identifies the states in the weakly localized regions for the real-world data. However, distinguishing between delocalized weakly localized and strongly localized states associated with real-world graphs poses a significant challenge. It might be due to the imbalance of data points in different states and limited dataset availability. Further, alongside the IPR value prediction task, we have also incorporated Shannon entropy (SE) as a measure of localization, which is predicted using a GCN model (Appendix F). One can use Rényi entropy instead of SE, and this direction will be explored in the future. This work focuses solely on the principal eigenvector and provides a first proof of concept that GNN models can identify localized and delocalized states in linear dynamics on binary undirected graphs. The models can also be readily extended to study the localization behavior of other eigenvectors and to directed or weighted graphs [35].
Code availability: The code used in our experiments is available in the GitHub repository at https://github.com/amitrezaiwf/Graph-Neural-Network-For-Steady-State-Identification, enabling reproducibility and facilitating future research built upon this work.